OpenStudy (anonymous):

Solve the equation 19y - 3(y-6) = 66

OpenStudy (callisto):

19y - 3(y-6) = 66 19y -3y +18 =66 16y = 48 y=3

OpenStudy (anonymous):

19y-3(y-6)=66 19y-3y+18 = 66 16y=66-18 16y=48 y=3